OpenStudy (anonymous):

24 is 15% of what number?

OpenStudy (anonymous):

Please use this for any of those problems p/100 = is/of. So, for your problem we'll get 15/100=24/x. Cross multiply and solve for x. Ok?

OpenStudy (anonymous):

Let the number be: x \[\frac{15}{100}x = 24\] \[x = \frac{24 \times 100}{15}\] Solve this to get your answer..